# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v3

@@ IV MAP Analytical Bound for GT with General Correlated Prior Statistics
In this section, we provide a sufficiency bound for the number of required pooled tests for any MAP decoder with an arbitrary underlying correlation between infected items. This bound also applies to the LVA-based MSGT algorithm, which can approach MAP decoder performance for a sufficiently large list [32]. We calculate this bound by extending the known result of maximum likelihood (ML) decoding for Group Testing (GT) problems [17], which bounds the probability of error of a ML decoder using a Gallager-type bound [44] for multiple access channel (MAC) channels [50]. Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword. While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [51, 52, 53, 54, 55, 56, 57], they do not expand naturally to GT, since GT can be thought as a special case where only “$K$” users transmit at any given point. This key difference necessitates dedicated analysis as follows in this section. In particular, to gain from the correlated prior statistics information, unlike ML decoders and bounds e.g., in [17, 14]. We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5.
@@ IV-A Definitions and notations
Let $\left[n\right]=\left\{1,\ldots,n\right\}$. Let $H(\cdot)$ and $I(\cdot;\cdot)$ denote Shannon’s entropy and mutual information, respectively [58]. All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise. Let $\omega\in\left[\binom{N}{K}\right]$ denote the index of a specific set of $K$ items out of the $N$ items. Throughout the paper, $\omega^{*}$ denotes the index of the set of the $K$ actual infected items. Let $W$ denote the random variable which points to the index of the infected items, and has a probability distribution of $P_{W}\left(\cdot\right):\left[\binom{N}{K}\right]\rightarrow\left[0,1\right]$. In the following, $W$ might be omitted for convenience of read. The $\omega$-th set is denoted by $S_{\omega}$. Note that $\lvert S_{\omega}\rvert=K$. The rows of $\mathbf{X}$ that correspond to the items in $S_{\omega}$ are denoted by $\mathbf{X}_{S_{\omega}}$. The output For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$. Note that $\lvert FN\left(\omega,\omega^{\prime}\right)\rvert=\lvert FP\left(\omega,\omega^{\prime}\right)\rvert$. In the following, we omit the dependency on $\omega,\omega^{*}$ and write $\mathbf{X}_{TP},\mathbf{X}_{FP},\mathbf{X}_{FN}$ instead of $\mathbf{X}_{TP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FN\left(\omega,\omega^{*}\right)}$ for notational simplicity. We say that an error of size $i$ happens if an output set of items $\omega\neq\omega^{*}$ is chosen and $\lvert FP\left(\omega,\omega^{*}\right)\rvert=i$. This event is denoted by $E_{i}$. Assuming the event $E_{i}$ happened, the index of the FN items is denoted by $j$, i.e. $j\in\left[\binom{N-K}{i}\right]$. Let $\Omega$ denote a function that converts a $j\in\left[\binom{N-K}{i}\right]$ index to a set of size $i$ of the FN items (the dependence on $\omega^{*}$ is omitted for notational simplicity). A decoder is denoted by $g$ and outputs some index $\omega=g\left(Y^{T}\right)$, where the dependency on $\mathbf{X}$ is omitted in favor of notational simplicity. These definitions are illustrated in the following example.
@@ IV-C Sufficiency GT MAP Bound
$\begin{split}&E_{0}(\rho)=\\ &-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
@@
for $0\leq\rho\leq 1$.
@@
$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)$
-- $\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})$
$\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}=I(X_{FP};X_{TP},Y).$
@@
This is obtained directly from [59] as also elaborated in [57, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
$T\emph{f}(\rho)=$ $T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)-\frac{H\left(P_{S_{2}|S_{1}}\right)}{T}\Bigg{)}$
-- $\log\frac{1}{\binom{K}{i}}+T\frac{\rho^{2}}{2}E_{o}^{\prime\prime}\left(\psi\right)-T\frac{\rho^{2}}{2}E_{s,j}^{\prime\prime}\left(\psi\right),$
@@
$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}$
$+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
@@
Claim 1 ([17] and [45]).
For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by $I(X_{FP};X_{TP},Y)\geq\frac{i}{K}$.
@@ V Numerical Evaluation
This section assesses the performance of the proposed MSGT algorithm by numerical study. First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III. Then, in Subsection V-2, we contrast the performance of MSGT with the one of DD, ML, and MAP in a practical regime of $N$ and $K$. To generate the correlated prior information between adjacent items, we use the Gilbert-Elliot (GE) model [60] (also considered very recently analytically in [61]), which is a stationary Markov chain of order 1 with two states: one representing an error phase and the other an error-free phase. The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$. These characteristics align well with the inputs required by the MSGT algorithm. In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [34, 35], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 62]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
@@ V-1 Theoretical Analysis
$T\geq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right).$
This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5.
@@ V-2 Algorithm Evaluation
We demonstrate the performance of MSGT using simulation. The population is sampled from GE model, and the regime is $K=\mathcal{O}\left(N^{\theta}\right)$ with $\theta\leq 1/3$. The GE parameters serve as our prior statistics, but in practice, we ignore samples where the number of defective items does not match $K$. In addition, although in the complexity analysis, we considered Bernoulli encoder for simplification, here we use a near-constant column weight encoder that optimizes DND’s performance [37], with $\frac{\ln{2}}{K}T$ tests sampled randomly for each item. The parameter $\gamma$ was chosen to satisfy (3) and $L=500$ was chosen empirically. Fig. 5 compares MSGT to MAP, ML, and DD algorithms. We run DND and DD before ML and MAP for reasonable runtime and memory consumption. The population includes $N\in\{500,1024,10000\}$ items and $K\in\{3,8,13\}$ defective items, respectively, and the empirical success probability is the average over $1000$ experiments. Note that for $N=1024,10000$ it is no longer possible to compare the performance since ML and MAP become infeasible (Fig. 5(b),5(c)).
@@
VI Discussion and Future Work
To the best of our knowledge, the existing solutions in the literature do not offer an efficient and general solution for the regime below $T_{ML}$. In this work, we focus on this regime of the number of tests (i.e., $T<T_{ML}$), where the proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of the potential defective combinations. This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.

### v4

@@ IV MAP Analytical Bound for GT with General Correlated Prior Statistics
In this section, we provide a sufficiency bound for the number of required pooled tests for any MAP decoder with an arbitrary underlying correlation between infected items. This bound also applies to the LVA-based MSGT algorithm, which can approach MAP decoder performance for a sufficiently large list [32]. We calculate this bound by extending the known result of ML decoding for GT problems [17], which bounds the probability of error of a ML decoder using a Gallager-type bound [53] for multiple access channel (MAC) channels [59]. Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword. While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [36, 37, 38, 39, 40, 41, 42], they do not expand naturally to GT, since GT can be thought as a special case where only “$K$” users transmit at any given point. This key difference necessitates dedicated analysis as follows in this section. In particular, to gain from the correlated prior statistics information, unlike ML decoders and bounds e.g., in [17, 14]. We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5. Finally, in Subsection IV-D, we evaluate this bound in Gilbert-Elliott (GE) model [60] and illustrate it for a practical regime in Section V.
@@ IV-A Definitions and notations
Let $\left[n\right]=\left\{1,\ldots,n\right\}$. Let $H(\cdot)$ and $I(\cdot;\cdot)$ denote Shannon’s entropy and mutual information, respectively [61]. Let $\omega\in\left[\binom{N}{K}\right]$ denote the index of a specific set of $K$ items out of the $N$ items. Throughout the paper, $\omega^{*}$ denotes the index of the set of the $K$ actual infected items. Let $W$ denote the random variable which points to the index of the infected items, and has a probability distribution of $P_{W}\left(\cdot\right):\left[\binom{N}{K}\right]\rightarrow\left[0,1\right]$. In the following, $W$ might be omitted for convenience of read. The $\omega$-th set is denoted by $S_{\omega}$. Note that $\lvert S_{\omega}\rvert=K$. The rows of $\mathbf{X}$ that correspond to the items in $S_{\omega}$ are denoted by $\mathbf{X}_{S_{\omega}}$. The output For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$. Note that $\lvert FN\left(\omega,\omega^{\prime}\right)\rvert=\lvert FP\left(\omega,\omega^{\prime}\right)\rvert$. In the following, we omit the dependency on $\omega,\omega^{*}$ and write $\mathbf{X}_{TP},\mathbf{X}_{FP},\mathbf{X}_{FN}$ instead of $\mathbf{X}_{TP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FN\left(\omega,\omega^{*}\right)}$ for notational simplicity. We say that an error of size $i$ happens if an output set of items $\omega\neq\omega^{*}$ is chosen and $\lvert FP\left(\omega,\omega^{*}\right)\rvert=i$. This event is denoted by $E_{i}$. Assuming the event $E_{i}$ happened, the index of the FN items is denoted by $j$, i.e. $j\in\left[\binom{N-K}{i}\right]$. Let $\Omega$ denote a function that converts a $j\in\left[\binom{N-K}{i}\right]$ index to a set of size $i$ of the FN items (the dependence on $\omega^{*}$ is omitted for notational simplicity). A decoder is denoted by $g$ and outputs some index $\omega=g\left(Y^{T}\right)$, where the dependency on $\mathbf{X}$ is omitted in favor of notational simplicity. These definitions are illustrated in the following example.
@@ IV-C Sufficiency GT MAP Bound
$\begin{split}&E_{0}(\rho)=-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
@@
for $0\leq\rho\leq 1$, where $\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\}$, $\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\}$, and $S_{\omega}=S_{1}\cup S_{2}$.
@@
$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)-\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}$
$=$ $I(X_{FP};X_{TP},Y).$
@@
This is obtained directly from [62] as also elaborated in [42, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
$T\emph{f}(\rho)=$ $T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)-\frac{H\left(P_{S_{2}|S_{1}}\right)}{T}\Bigg{)}-\log\frac{1}{\binom{K}{i}}+T\frac{\rho^{2}}{2}E_{o}^{\prime\prime}\left(\psi\right)-T\frac{\rho^{2}}{2}E_{s,j}^{\prime\prime}\left(\psi\right),$
@@
$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
@@
Claim 1 ([17]).
For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by
$I(X_{FP};X_{TP},Y)\geq\frac{i}{K}.$
@@
(c) Minimum $\gamma$ parameter to satisfies (3) for $N=500$, and $K=3$.
IV-D Gilbert Elliott calculations
Consider a GE model [60]. It is a Markov process with two states. In the context of GT, as also recently considered in [63], we call them infected (state 1) and not infected (state 0). The probability of transition from state 0,1 to state 1,0 is denoted by $q,s$, respectively. The probability that the first state is in state 1 is denoted by $\pi\triangleq P\left(U_{i}=1\right)$. Unless stated otherwise, we assume the system is in steady state, that is, the probability of the system to be at state 1 at any point is given by $\pi=\frac{q}{q+s}$.
To evaluate Theorem 5, we need to evaluate
$H\left(P_{S_{2}|S_{1}}\right)={\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)\log\frac{P\left(S_{1}\right)}{P_{W}\left(S_{1},S_{2}\right)}\right),$
where $P\left(S_{1}\right)=\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)$.
Let $S_{\omega}=\left(S_{1},S_{2}\right)$ denote a specific set of $K$ items. Recall that $P_{W}\left(S_{\omega}\right)$ calculates the probability to get $S_{\omega}$ as the infected set, given that there are $K$ infected items.
Let $\mathcal{G}\mathcal{E}_{N}$ denote a GE process that runs for $N$ steps. Let $\left\{U_{i}\right\}_{i=1}^{N}$ denote the states of the process during the $N$ steps. Let $\lvert\mathcal{G}\mathcal{E}_{N}\rvert\triangleq\sum_{i=1}^{N}U_{i}$. While our analysis is conditioned on the assumption that exactly $K$ items are infected, a general $\mathcal{G}\mathcal{E}_{N}$ process may result in any number of infected items between 0 and $N$. We therefore analyze and simulate a special class of GE processes; GE processes that reach the infected state exactly $K$ times after $N$ steps. Such a GE process that is guaranteed to produce exactly $K$ infected items is denoted by $\mathcal{G}\mathcal{E}_{N,K}$. Then $P_{W}\left(S_{\omega}\right)$ of $\mathcal{G}\mathcal{E}_{N,K}$ is given as follows
$P_{W}\left(S_{\omega}\right)$ $=$ $P\left(\mathcal{G}\mathcal{E}_{N,K}\text{ results in }S_{\omega}\right)$ (12)
=$\overset{(a)}{=}$ $P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\lvert\;\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)$
=$\overset{(b)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega},\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)}$
=$\overset{(c)}{=}$ $\frac{P\left(\mathcal{G}\mathcal{E}_{N}\text{ results in }S_{\omega}\right)}{P\left(\lvert\mathcal{G}\mathcal{E}_{N}\rvert=K\right)},$
where (a) follows from the definition of $\mathcal{G}\mathcal{E}_{N,K}$, (b) from the definition of conditional probability, and (c) follows since $S_{\omega}$ contains exactly $K$ infected items.
IV-D1 Numerator of Eq.(12)
The numerator evaluates the probability that a standard GE process results in a specific pattern of states, where state one is achieved only at the states in $S_{\omega}$. Recall that GE is a Markov chain, and so $P\left(S_{\omega}\right)=P\left(U_{1}\right)\cdot\prod_{i=2}^{N}P\left(U_{i}\lvert U_{i-1}\right)$. This can be readily calculated using the definitions of $q,s,\pi$, as described at the beginning of this section
$P\left(U_{1}\right)=\begin{cases}1-\pi&\text{if }1\not\in S_{\omega}\\ \pi&\text{if }1\in S_{\omega}\end{cases},$
and
$P\left(U_{i}\lvert U_{i-1}\right)=\begin{cases}1-q&\text{if }i\not\in S_{\omega},i-1\not\in S_{\omega}\\ q&\text{if }i\in S_{\omega},i-1\not\in S_{\omega}\\ s&\text{if }i\not\in S_{\omega},i-1\in S_{\omega}\\ 1-s&\text{if }i\in S_{\omega},i-1\in S_{\omega}\end{cases}.$
IV-D2 Denominator of Eq.(12)
We create a new Markov chain that represents $\sum_{i=1}^{n}U_{i}$ as $n$ advances from $1$ to $N$. This Markov chain keeps track of how many times the GE process was in state 1. Hence this Markov chain has $2\left(N+1\right)$ states. State $n,n^{\prime}\in\left\{0,\ldots,N\right\}$ denotes a state where the GE process was in the infected state (state 1) $n$ times, and the current state is 0,1, respectively. An example of this Markov chain is shown in Fig. 6. For convenience, the states $0^{\prime},N$ are also listed, even though they cannot be achieved. This Markov chain’s states can be arranged in a vector, where the $j$-th state represents $\sum_{i=1}^{N}U_{i}=\left\lfloor j/2\right\rfloor$, and even/odd $j$ represents the current state is not/is 00. For instance, in the example of Fig. 6, the states $6/7$ represent states 2/$2/2^{\prime}$ in the figure. The initial state vector is given as by $\left(1-\pi,0,0,\pi,0,\ldots,0\right)$. The transition matrix is given as follows
$P=\begin{bmatrix}1-q&0&0&q&0&0&\dots&0\\ 0&0&0&0&0&0&\dots&0\\ 0&0&1-q&0&0&q&\dots&0\\ 0&0&s&0&1-s&0&\dots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&0&0&\dots&0\end{bmatrix}.$
(a) $N=500$ and $K=3$.
(b) $N=1024$ and $K=8$.
(c) $N=10000$ and $K=13$.
The probabilities to be in any state after $N$ steps is therefore given by
$\left(1-\pi,0,0,\pi,0,\ldots,0\right)\cdot P^{N-1}.$ (13)
To conclude the calculation, the probability to have exactly $K$ infected items after $N$ steps can be obtained by summing items $12K+$ 1 and $22K+$ 2 of the vector in (13).
Figure 6: Markov chain for the number of infected items of a GE model, $N=4$.
In the next section, we evaluate the proposed sufficient bound given in Theorem 5 for MAP decoder in the GE setting considered herein.
@@ V Numerical Evaluation
This section assesses the performance of the proposed MSGT algorithm by numerical study. First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III. Then, in Subsection V-2, we contrast the performance of MSGT with those of DD, ML, and MAP in a practical regime of $N$ and $K$. To generate the correlated prior information between adjacent items, we use the special GE process that reaches the infected state exactly $K$ times after $N$ steps, as described and analyzed in Section IV-D. The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$. These characteristics align well with the inputs required by the MSGT algorithm. In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [43, 44], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 64]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
(a) Number of computational operations.
(b) Correlation matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
(c) Long memory priors - exact recovery.
@@ V-1 Theoretical Analysis
$T\geq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right).$ (14)
This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5. In Fig 8, we compare the lower and the upper bounds with GE model. The lower and upper bounds for GT with correlated prior statistics are given in (14) and in Section IV-D, respectively, while for GT without prior statistical information, the bounds are as given in [17].
Figure 8: Lower and upper bounds for GT with GE model (left). MSGT and corresponding MAP upper bound for GT with GE model (right). For MAP and ML upper bounds $\varepsilon=0.6$.
@@ V-2 Algorithm Evaluation
We demonstrate the performance of MSGT using simulation. The population is sampled from GE model, and the regime is $K=\mathcal{O}\left(N^{\theta}\right)$ with $\theta\leq 1/3$. The GE parameters serve as our prior statistics, but in practice, we ignore samples where the number of defective items does not match $K$. In addition, although in the complexity analysis, we considered Bernoulli encoder for simplification, here we use a near-constant column weight encoder that optimizes DND’s performance [46], with $\frac{\ln{2}}{K}T$ tests sampled randomly for each item. The parameter $\gamma$ was chosen to satisfy (3) and $L=500$ was chosen empirically. Fig. 5 compares MSGT to MAP, ML, and DD algorithms. We run DND and DD before ML and MAP for feasible runtime and memory consumption. The population includes $N\in\{500,1024,10000\}$ items and $K\in\{3,8,13\}$ defective items, respectively, and the empirical success probability is the average over $1000$ experiments. Note that for $N=1024,10000$ it is no longer possible to compare the performance since ML and MAP become infeasible (Fig. 5(b),5(c)).
@@
VI Conclusions and Future Work
In this work, we propose a MAP decoder for GT problems with arbitrary correlated sources. We analyze the performance of the MAP decoders for this case and provide, to the best of our knowledge, the only analysis that goes below $T_{ML}$, as the existing solutions in the literature do not offer an efficient and general solution for this regime. Our proposed MSGT can leverage the prior correlated information, within the LVA step, to significantly reduce the number of potential defective combinations. This approach yields an efficient computational solution that, as demonstrated in our simulation result for practical scenarios, can approach the minimum number of tests as MAP algorithm.
Future work includes the study of the proposed algorithm and analytical bounds in communication systems [36, 37, 38, 39, 40, 41, 42] and signal processing applications [23, 13, 24], in which correlation between different frequencies, time signals, or among different sensors, etc., can also be utilized to achieve more precise estimations.